# VisFly: An Efficient and Versatile Simulator for Training Vision-based Flight

### v1

@@ III-B Scenes
We use Habitat-sim as the engine to render the visual observation of AI agents. Rapid rendering is one of the most crucial character of habitat-sim. We evaluate the frame rate of VisFly while running with physics, and show the result in Figure 3. With visual input, VisFly can reach the frame rate of up to 10000 FPS, almost 100 times of those using Unity or Unreal Engine.
Figure 3: Frame rate of VisFly. We evaluate the frame rate using a desktop with RTX 4090 and 32-core 13th Gen Intel(R) Core(TM) i9-13900K. The multiple agents run repeatedly with random actions. Left sub-figure is the result tested in Replica datasets while total number of agents is 100. The maximum frame rate can reach as high as 10,000 FPS. When the resolution is set to 256$\times$256, a popular resolution choice in open-source models and hardware, the frame rate can also reach up to 6,000 FPS. The right sub-figure illustrates the trend of how the frame rate varies with the number of scenes running simultaneously when resolution is 64$\times$64. Increasing number of scenes decreases the simulating speed while the degradation can be tolerated.
Meanwhile, a series of datasets including HM3D[32], MP3D[33], Replica[34], and HSSD[35] are compatible with habitat-sim. In these datasets, some of the data scanned from the real world are highly photo-realistic. Users can directly utilize these open-source datasets to evaluate or train models. Except for open datasets, we also define datasets generator to generate basic cluttered environments for common usage like training navigation policy. These real-world scenarios, combined with powerful scene customizable features, make VisFly more realistic, effectively reducing the adaptation challenges for robots during Sim2Real transitions. Rich extant datasets eliminate the extra efforts for gathering the original data, and offer a fairer approach compared to self-defined environments and promising to serve as baselines for quadrotor tasks. VisFly supports users to equip quadrotors with several different types of sensors, reserving extension for developing complicated multi-camera aerial robots.
@@
Specialized for ground robots, habitat-sim assumes that robots always operate on a 2-dimensional plane. It precomputes the NavMesh, which projects spatial objects onto the floor and generates distance information to the closest obstacle. This approach is not suitable for aerial robots that move freely. Therefore, we have developed a spatial distance computation module for the modified habitat-sim, enabling it to detect the relative position of the nearest collision point. The modified version of habitat-sim will also be released within the same website.
@@
VisFly supports parallelized agents as well as multi-agents in one scene. These two modes share the same underlying framework. The key difference in outcomes is that the former visually omits other drones and ignores collisions between them, whereas the latter does the opposite.
III-C Sim-to-Real Transfer
@@
The gap between simulation and reality hinders the usage transfer of robots trained in simulation. In order to upgrade the performance of practical implementation, one effective approach is to improve the realism of simulation, whose drawback is discussed in Introduction. The other approach is to enlarge the data distribution generated by simulator, strengthening the in-distribution generalization of model, which is called as domain randomization[7]. Considering the real flight scenarios, we add customizable noise to all the sensors as Table II.
@@
The initial position, velocity, orientation, angular velocity, can also be randomized to follow a normal or uniform distribution. To avoid unstable training, agents are reset with initial position away from obstacles in cluttered environments. An important contribution of VisFly is that multiple scenes are able to run distributively and return observations centrally together. During training, scenes are treated as a datasets, that means, scenes could be sampled and applied with shuffle option. To the best of our knowledge, VisFly is the first simulator that runs multiple scenes simultaneously to achieve scene randomization.
III-D Gym-wrapped Environment
Gym is an open-source library of several regular interactive environments designed for reinforcement learning. It wraps the processes of environment and set interfaces for user to interact with them. Users are able to obtain the agents state including reward, observation, finished status, running information through step() function.
VisFly’s ultimate mission is to facilitate the development of aerial robots learning for several tasks. Therefore, we have taken the extensibility of the basic environment into account as much as possible when packaging the environment. Two abstract environment class droneGymEnvsBase, multiDroneGymEnvsBase are provided. When users design custom tasks for a quadrotor, they need only define a class that inherits from the abstract gym environment class and override three functions: get_success(), get_reward(), and get_observation() as shown in Algorithm 1. This approach creates a new gym environment tailored to the specific mission. We provide three examples in Section IV for users to reference. This customizable architecture enables the training of models that can act as backbones for more complex tasks. For instance, a model trained for obstacle avoidance can serve as the baseline or backbone for higher-level autonomous tasks, such as search and rescue operations in wilderness areas.
Algorithm 1 High-level Tasks Definition
class DroneGymEnvsBase
         @abstractmethod
         def get_reward()
         @abstractmethod
         def get_observation()
         @abstractmethod
         def get_success()
class DroneNavigationEnvs(DroneGymEnvsBase)
         def get_reward()
               return $r_{distance}+r_{speed}-r_{collision}$
         def get_observation()
               return $state+vision+target$
         def get_success()
               return $status_{reach\_the\_target}$
class DroneLandingEnvs(DroneGymEnvsBase)
         def get_reward()
               return $-r_{height}+r_{speed}-r_{collision}$
         def get_observation()
               return $state+vision+visual\_target$
         def get_success()
               return $status_{height<=0.1}\&status_{speed<=0.1}$
IV Applications
In order to verify this simulator, we list three instances: (1) learning to land, (2) learning to navigate in cluttered environment, (3) learning to cooperatively cross the narrow gap. We choose reinforcement learning to train an end2end model since RL is data-hungry. In the following instances, considering that the demonstration point is relative with training efficiency instead of algorithm categories, we only employ PPO [37] from stable-baselines3[38] (SB3) for its outstanding capacity to handle high-dimensional observation, to train the models. The backbone architecture of the policy network is consistent with SB3 as Figure 4, and we modify it to make it more accessible for any tasks.
@@ IV Applications
The image extractor is a conventional neural network that comprises three layers, each containing a convolutional layer, an activation function, a max pooling operation, and concludes with a linear layer with 128 units. The convolutional layers have kernel sizes of 5, 3, and 3, with corresponding channel counts of 6, 12, and 18. Each max pooling operation uses a kernel size of 2. The target extractor, state extractor, swarm extractor are all multilayer perceptrons (MLPs) that comprise two linear layers with unit counts of 128, 64. Both action net and value net are also MLPs that comprise two linear layers with unit counts of 64, 64. All the activation function are ReLU. Noting that, this architecture has been parameterized, thus making it convenient to fine-tune. The action types in all the instances are CTBR. And the quadrotor take its full state (position, orientation, linear velocity, angular velocity), visual image, target as observation. Most of training hyper-parameters are set as default, detailed setting, as well as reward function, done function, observation function, are introduced in source code.
@@
Figure 4: Neural network architecture illustration of policies. The image feature extractor has been modified to be more customizable, including backbones such as ResNet[39], MobileNet[40], EfficientNet[41]. The policies for three tasks follows this architecture but slightly different. VisFly additionally retains the recurrent network interface [42] for training tasks that do not strictly conform to Markov process.
@@ IV-A Learning to Land
Quadrotor begins from aerial flight status, decelerates, and eventually lands stably on the ground. It continuously monitors the limited landing area directly below. This landing area is a black rectangle positioned horizontally on the ground within a clear garage, as shown in Figure 5. Quadrotor calculates the center of the $0.5\times 0.5m^{2}$ black rectangle in their field of view, strives to keep it centered, and incorporates it as part of their observations. During flight, quadrotor only acquires $64\times 64$ RGB image from camera instead of any precise spatial location of landing area. The center of landing area is calculated using threshold segmentation and incorporated as target observation. This training takes 1.7 hour, and the additional training time from the Section IV-B task accounts for the center preprocessing required for the landing area.
@@
IV-B Learning to Navigate in Cluttered Environment
The quadrotor should reach the target position safely while avoiding the obstacles during flight. It is one of the most fundamental tasks for quadrotors in real-world applications. We set a garage scene filled with random stones of density of $0.15/m^{3}$ in the spatial space as shown in Figure 5. It is more challenging more than dense scenes of trees or columns, since the quadrotor considers the vertial motion as well. Quadrotor starts from initial position, flies across the stone region and finally pass the target position. It takes $64\times 64$ depth image as visual observation and final position as target observation. This training takes 0.73 hour. To the best of our known, this instance is possible to be the first end2end autonomous navigation in cluttered environment without any training tricks, curriculum, privileged information or customized primitives within simulation. It is noted that, not the elegant reward setting but the abundant data contributes to the training.
IV-C Cooperatively Learning to Cross the Gap
We inserted two walls in the clear garage to create a narrow gap between them as Figure 5. Three quadrotors start from one side of the wall, fly towards distributed targets in the other side, and cross the narrow gap simultaneously. They must cooperate by considering each other’s state, to safely avoid both obstacle and other agents. Quadrotors take $64\times 64$ depth image as visual observation, final position as target observation, additionally other agents’ state as swarm observation. Unlike single-agent tasks, VisFly allows only one set of swarm in one scene, which limits the maximum batch size due to memory constraints. This training takes 1.1 hour. The result shows that, the by observing the other agents’ state, quadrotors are able to avoid collision with obstacles as well as each other.
@@ V Discussion
VisFly’s torch-based dynamics achieve differentiable physics at the expense of lower accuracy compared to professional simulators like Gazebo. However, this trade-off does not negatively impact Sim2Real, even in extreme scenarios such as quadrotor racing[27], where the quadrotor must fly close to its dynamic upper bound. Additionally, VisFly does not integrate high-level controllers such as model predictive controller (MPC) for two primary reasons. Firstly, these advanced controllers consume excessive computational resources in both simulation and real flight. Secondly, end2end learning is preferred as it more closely imitates natural decision-making processes. In future work, we will incorporate real-world dynamics into VisFly to make HITL available.
@@ VI Conclusion
In this paper, we introduce a both photo-realistic and sampling-efficient quadrotor simulator VisFly. VisFly integrates differentiable dynamics with habitat-sim, solving the shortage of visual training data especially for reinforcement learning. It wraps whole environments as gym interfaces, making it available to directly implement the latest reinforcement learning algorithms, such as stable-baselines3. It offers parallelized simulation to perform multi-agent system as well as multi-scene system. Since habitat-sim is widely used, it can also leverage a large amount of open-source scene datasets specifically designed for embodied learning, without pre-processing. We conducted three learning tasks involving visual input: learning to land, learning to navigate through a cluttered environment, and swarm learning to cross a narrow gap. The results confirm the reliability of this framework and provide foundational instances that enable users to extend high-level tasks. According to our current understanding, it could be the first quadrotor simulator that intensively generates available visual data in a short time. It is hopeful that VisFly could not only bridge the gap between simulation and reality but also between embodied learning and aerial robotics.
Acknowledgments
This work is inspired by Habitat-sim[20]. Part of the code architecture refers to Flightmare[17]. We sincerely appreciate their contributions to the quadrotor and embodied learning communities.

### v3

@@ III-B Efficient rendering and open-source scene management
We use the rapid rendering engine of Habitat-sim to render the visual observation data efficiently. Figure 3 presents the frame rate of VisFly while running physics simulations. With visual input, VisFly can achieve a frame rate of up to 10,000 FPS, almost 100 times faster than simulators using Unity or Unreal Engine.
Figure 3: Frame rate performance of VisFly. Left: Tested with 100 agents in the Replica dataset, achieving up to 10,000 FPS. At 256$\times$256 resolution, the frame rate still reaches up to 6,000 FPS. Right: Variation in frame rate with the number of scenes running simultaneously at 64$\times$64 resolution.
Similar to Habitat-Sim, VisFly supports a range of open-source 3D scene datasets, including HM3D[33], MP3D[34], Replica[35], and HSSD[36], many featuring highly photorealistic real-world scans. These datasets facilitate direct training and evaluation of models. To augment these resources, we incorporate a scene generator for producing random, cluttered environments, such as those used in collision avoidance training. The combination of real-world data and customizable scenes enhances VisFly’s adaptability, reducing Sim-to-Real challenges. By eliminating the need for extensive data collection, these datasets also foster fair comparisons and establish strong baselines for quadrotor research. VisFly supports common vision and motion sensors, and is extensible to complex multi-camera systems.
@@
One requirement in training flight policies is to query the nearby obstacles on the fly. Habitat-sim, specialized for ground robots, assumes that robots operate on a 2-dimensional surface. It precomputes the NavMesh, which projects spatial objects onto the floor and generates distance information to the closest obstacle. This approach is not suitable for aerial robots that move in 3D space. Therefore, we developed a spatial distance computation module via the CGAL [^5] https://www.cgal.org/ package in C++, enabling it to detect the relative position of the nearest point on the obstacle without collision actually happening.
@@
VisFly supports parallelized agents as well as multi-agents within the same scene. These two modes share the same underlying framework. The key difference between them is that parallelized agents visually omit other drones and ignore collisions between them, while multi-agent scenarios include interactions and collisions among all drones.
III-C Modular design for Gym integration
Gym [24] is an open-source toolkit providing a standard interface for reinforcement learning. It offers a collection of pre-built environments and a framework for custom environment development. Agents are able to interact with environments through step() function, obtaining rewards, observations, termination signals, and other information at each time step.
VisFly’s ultimate mission is to facilitate the development of aerial robot learning for autonomous flight tasks. Therefore, we have taken the extensibility of the basic environment into account as much as possible when packaging the environment. Two abstract environment classes droneGymEnvsBase, multiDroneGymEnvsBase are provided. When users design custom tasks for a quadrotor, they need only define a class that inherits from the abstract gym environment class and overrides three functions: get_success(), get_reward(), and get_observation() as shown in Lst. 1. This approach creates a new gym environment tailored to the specific mission. We provide three examples in Section IV for users to reference. This hierarchical architecture enables the training of models that can act as backbones for more complex tasks. For instance, a model trained for obstacle avoidance can serve as the low-level policy or backbone for higher-level autonomous tasks, such as search and rescue operations in wilderness areas.
1class DroneGymEnvsBase:
2 @abstract method
3 def get_reward()
4 @abstract method
5 def get_observation()
6 @abstract method
7 def get_sucess()
8 ...
9class DroneNavigationEnvs(DroneGymEnvsBase)
10 def get_reward()
11 ...
12 return r_distance+r_speed-r_collision
13 def get_observation()
14 ...
15 return state+vision+target
16 def get_sucess()
17 ...
18 return status_reach_the_target
19 ...
20class DroneLandingEnvs(DroneGymEnvsBase)
21 def get_reward()
22 ...
23 return -r_height+r_speed-r_collision
24 def get_observation()
25 ...
26 return state+vision+visual_target
27 def get_sucess()
28 ...
29 return status_height<0.1&status_speed<0.1
Lst. 1: Mocular class design for different tasks
III-D Domain randomization for sensor data and flight states
@@
The gap between simulation and reality usually hinders the real-world deployment of models trained in simulation. In order to reduce the sim-to-real domain gap, one effective approach is to improve the realism of simulation, whose drawback is discussed in Introduction. The other approach is to enlarge the data distribution generated by simulators, strengthening the in-distribution generalization of models, which is called domain randomization[15]. For training vision-based flight policies, it is important to add noises to both the motion and vision sensors as listed in Table II.
@@
In addition to adding noises to sensors, VisFly allows for the randomization of initial conditions, including position, velocity, orientation, and angular velocity, which can follow either a normal or uniform distribution. To achieve better generalization of the policy, agents are reset with initial positions away from obstacles in cluttered environments. An important contribution of VisFly is its capability to run multiple scenes distributively and aggregate observations centrally. During training, scenes are treated as datasets, allowing them to be sampled and shuffled. To the best of our knowledge, VisFly is the first simulator that supports simultaneous execution of multiple scenes for enhanced scene randomization.
IV Training examples
To validate this simulator, we present three cases of training vision-based flight policies: (1) learning to land, (2) learning to navigate in a cluttered environment, and (3) learning to cross the narrow gap cooperatively. We use reinforcement learning to train end-to-end models for controlling the drones using the vision inputs. To focus on training efficiency rather than the algorithms themselves, we employ PPO [38] from stable-baselines3[39] (SB3) for its outstanding capability to handle high-dimensional observationsThe backbone architecture of the policy network follows the standard design in SB3 , as shown in Figure 4, with modifications to enhance its adaptability across various tasks.
@@ IV Training examples
The image feature extractor is a conventional neural network that comprises three layers. Each layer includes a convolutional layer, an activation function, and a max pooling operation, followed by a linear layer with 128 units. The convolutional layers have kernel sizes of 5, 3, and 3, with channel counts of 6, 12, and 18. Each max pooling operation uses a kernel size of 2. The target extractor, state extractor, and swarm extractor are all multilayer perceptrons (MLPs) with two linear layers, having 128 and 64 units. The action net and value net are also MLPs, with two linear layers of 64 units. All activation functions are ReLU. This architecture has been parameterized, making it convenient to fine-tune. The action type in all the examples is CTBR. The quadrotor takes its full state (position, orientation, linear velocity, angular velocity), visual image, and target as observations. Most of the training hyper-parameters are set to default values. Detailed settings, the reward, done, and observation functions are provided in the source code.
@@
Figure 4: Neural network architecture used across different tasks. The image feature extractor has been modified to be more customizable, including backbones such as ResNet[40], MobileNet[41], EfficientNet[42]. The policies for the three tasks follow this architecture but are slightly different. VisFly additionally incorporates a recurrent network interface [43] for training tasks that do not strictly conform to the Markov process.
@@ IV-A Learning to land
In this task, the quadrotor starts in flight, gradually decelerates, and eventually lands safely on the landing area. It continuously monitors the limited landing area directly below, which is represented as a black rectangle positioned horizontally on the ground within a clear garage, as shown in Figure 5. The quadrotor calculates the center of the $0.5\times 0.5m^{2}$ black rectangle in its field of view, strives to keep it centered, and incorporates it into its observation. During the flight, the quadrotor captures only $64\times 64$ RGB image from the camera rather than acquiring precise spatial information about the landing area. The center of the landing area is determined using threshold segmentation and included as a target observation. This training process takes 1.7 hours, with additional time required for prepossessing the center of the landing area, as detailed in Section IV-B.
@@
IV-B Learning to navigate in a cluttered environment
In this task, the quadrotor must reach the target position safely while avoiding obstacles during flight. It is one of the most fundamental tasks for quadrotors in real-world applications. We set a garage scene filled with random stones of the density of $0.15/m^{3}$ in the spatial space as shown in Figure 5. This scenario is more challenging than dense scenes with trees or columns because it also requires the quadrotor to account for vertical motion. The quadrotor starts from an initial position, navigates through the stone-filled region, and eventually reaches the target position. It uses a $64\times 64$ depth image as the visual observation and the final position as the target observation. Training takes 0.73 hours. To our knowledge, this is the first instance of end-to-end autonomous navigation in a cluttered environment without employing training techniques such as curriculum learning, privileged information, or customized primitives within the simulation.
IV-C Learning to cross a narrow gap cooperatively
We inserted two walls in the clear garage to create a narrow gap between them as Figure 5. Three quadrotors start from one side of the walls, fly towards distributed targets on the other side, and pass through the narrow gap simultaneously. They must cooperate by considering each other’s state to avoid obstacles and other agents safely. Each quadrotor uses a $64\times 64$ depth image as its visual observation, the final position as its target observation, and the states of other agents as its swarm observation. Unlike single-agent tasks, VisFly supports only one swarm per scene, which limits the maximum batch size due to memory constraints. Training takes 1.1 hours. The results show that by observing the states of other agents and the environment, the quadrotors can effectively fly across the gap and avoid collisions with each other.
@@ V Discussion
VisFly’s PyTorch-based dynamics provide differentiable physics, though with slightly lower accuracy compared to professional simulators like Gazebo. However, lower accuracy does not impair the Sim-to-Real deployment. This has been demonstrated in challenging scenarios such as quadrotor racing [20] where the quadrotor is capable of flying close to its dynamic limits. Additionally, VisFly does not incorporate high-level controllers like Model Predictive Control (MPC) for two main reasons: First, these advanced controllers require substantial computational resources during simulation and real flight. Second, End-to-end learning is preferred because it closely mimics natural decision-making processes. In future work, we plan to integrate real-world dynamics into VisFly to enable Hardware-in-the-Loop (HITL) capabilities.
@@ VI Conclusion
This paper presents VisFly, a quadrotor simulator designed for efficient training of vision-based flight policies. Leveraging differentiable physics and Habitat-Sim’s rendering capabilities, VisFly supports single and multi-agent simulations across various controller types. Integrating with the gym environment provides a standardized interface for various reinforcement learning algorithms. Experimental results on landing, cluttered environment navigation, and cooperative gap crossing demonstrate the simulator’s effectiveness. VisFly serves as a robust foundation for future research, offering baseline implementations and extensibility for advanced quadrotor tasks.